Please help- I am so lost on this one!:
-2x+5x-9=3(x-4)-5
3x-9=3x+-12+-5
3x-9=3x+-17
+9 +9
3x=3x-8
3x/3=3x/3-8
x=x-8
+x+x
2x/2=-8/2 x=-4
so at what point am I going wrong? thanks!

Answer by stanbon(75887)

Please help- I am so lost on this one!:
-2x+5x-9=3(x-4)-5
3x-9=3x+-12+-5
3x-9=3x+-17
+9 +9
3x=3x-8
At this point if you subtract 3x from both sides you get:
0=-8
That is a contradiction which arose because you assumed all
the x's in the original problem could be replaced by the same
unknown number. It turned out that was not true and it led
to the contradiction.
So, your conclusion is "There is no solution"; the original
equation is inconsistent.
Cheers,
Stan H.